OpenStudy (anonymous):

The area of a rectangle is 21 ft^2 , and the length of the rectangle is 1ft less than twice the width. Find the dimensions of the rectangle.

OpenStudy (jdoe0001):

Area = length * width = \(\bf 21 ft^2\) w = width l = length length of the rectangle is 1ft less than twice the width -1 2 * w length = 2w -1  \(\bf Area = length \times width\implies 21 = l\times w\\ \quad \\ \implies 21 = (2w-1)\times w\)

OpenStudy (jdoe0001):

solve for "w" to see what the width is the length is 2w-1, so twice that much minus 1